OpenStudy (anonymous):

Solve using variation of parameters \[y''-2y'+y=e^{2x}\]

OpenStudy (turingtest):

complimentary....

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

\[y_c=c_1e^{-x}+c_2xe^{-x}\]

OpenStudy (turingtest):

right...

OpenStudy (turingtest):

now we need the wronskian, do you know about that?

OpenStudy (anonymous):

nope

OpenStudy (turingtest):

it is the determinant of a matrix

OpenStudy (turingtest):

for\[y_c=c_1y_1+c_2y_2\]the Wronskian is\[W(y_1,y_2)=\left|\begin{matrix}y_1&y_2\\y_1'&y_2'\end{matrix}\right|\]

OpenStudy (turingtest):

that's a determinant in case you forgot, so find that determinant

OpenStudy (anonymous):

\[\left|\begin{matrix}e^{-x}&xe^{-x}\\-e^{-x}&-e^{-x}(x-1)\end{matrix}\right|\] =\[e^{-x}(e^{-x}(x-1))-xe^{-x}e^{-x}\]

OpenStudy (anonymous):

completely wrong?

OpenStudy (anonymous):

oh it's plus

OpenStudy (turingtest):

yep you caught it :)

OpenStudy (anonymous):

\[e^{-x}(e^{-x}(x-1))+xe^{-x}e^{-x}\]

OpenStudy (anonymous):

that's it?

OpenStudy (turingtest):

you dropped a minus

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

\[-e^{-x}(e^{-x}(x-1))+xe^{-x}e^{-x}\]

OpenStudy (turingtest):

right

OpenStudy (turingtest):

simplify and be happy :)

OpenStudy (anonymous):

does it equal\[ e^{-x}\]? \[-e^{-x}(e^{-x}(x-1))+xe^{-x}e^{-x}=e^{-x}\] \[{-x}(e^{-x}(x-1))+xe^{-x}=1\]

OpenStudy (anonymous):

\[(e^{-x}(x-1))+xe^{-x}=1\]

OpenStudy (anonymous):

typo

OpenStudy (turingtest):

no, why did you set it equal to e^-x, we can't set it equal to anything yet...

OpenStudy (anonymous):

oh my bad haha

OpenStudy (anonymous):

so how should I simplify it then?

OpenStudy (turingtest):

\[-e^{-x}(e^{-x}(x-1))+xe^{-x}e^{-x}=e^{-x}(e^{-x}(1-x))+xe^{-x}e^{-x}\]now try again while I try to eat ;)

OpenStudy (anonymous):

-(x-1)=1-x 1-x=1-x

OpenStudy (anonymous):

no no no ...wait a second

OpenStudy (anonymous):

now i get 2-x=2-x

OpenStudy (anonymous):

nope still wrong

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

stop laughing

OpenStudy (anonymous):

Let's try that again \[-e^{-x}(e^{-x}(x-1))+xe^{-x}e^{-x}=e^{-x}(e^{-x}(1-x))+xe^{-x}e^{-x}\]

OpenStudy (anonymous):

they're just equal to each other

OpenStudy (anonymous):

\[-e^{-x}(e^{-x}(x-1))+xe^{-x}e^{-x}\]

OpenStudy (anonymous):

\[-e^{-2x}(x-1)+xe^{-2x}\]

OpenStudy (anonymous):

\[e^{-2x}((1-x)+x)\]

OpenStudy (anonymous):

\[e^{-2x}\]

OpenStudy (anonymous):

yes?

OpenStudy (turingtest):

yes, sorry, I was afk

OpenStudy (anonymous):

LOL ok

OpenStudy (turingtest):

do you know the formula for the particular solution for variation of parameters?

OpenStudy (anonymous):

something like that?

OpenStudy (turingtest):

almost unidentifiable to me in that form, but correct I'm sure I have the more direct formula memorized http://tutorial.math.lamar.edu/Classes/DE/VariationofParameters.aspx (the green box with the formula that says "variation of parameters)

OpenStudy (anonymous):

is the first one a y''? \[y''+q(t)y'+r(t)y=g(t)\]

OpenStudy (turingtest):

yeah, it looks screwy I know

OpenStudy (anonymous):

do we know what q(t) and r(t) are? are they the coefficients of y' and y? Probably not it looks like

OpenStudy (turingtest):

yes and in this case they are the coefficients, which in this case are constants the formula is just pointing out that they may not be

OpenStudy (anonymous):

oh I see, soo... \[y''+q(t)y'+r(t)y=g(t)\] \[y''-2y'+y=e^{2x}\] so now I integrate

OpenStudy (turingtest):

right, using the formula... r(t) and r(t) don't even come into the formula for the particular solution in variation of parameters

OpenStudy (turingtest):

q(t) and r(t)... they determine the complimentary only, which as you can see *is* part of the formula for the particular

OpenStudy (anonymous):

\[Y_p(t)=-e^{-x}\int\frac{xe^{-x}e^{-x}}{e^{-2x}(e^{-x},xe^{-x})}dx+xe^{-x}\int\frac{e^{-x}e^{2x}}{e^{-2x}(e^{-x},xe^{-x})}dx\]

OpenStudy (turingtest):

the first integrand is wrong I think

OpenStudy (turingtest):

I think you put the wrong g(x)

OpenStudy (anonymous):

\[Y_p(t)=-e^{-x}\int\frac{xe^{-x}e^{2x}}{e^{-2x}(e^{-x},xe^{-x})}dx+xe^{-x}\int\frac{e^{-x}e^{2x}}{e^{-2x}(e^{-x},xe^{-x})}dx\]

OpenStudy (turingtest):

oh wait, wait, the denominator in each integrand is the wronskian, which we found to be e^(-2x) I think you are confusing notation with the parentheses

OpenStudy (anonymous):

Oh lol so the denominator is just e^{-2x}

OpenStudy (turingtest):

right

OpenStudy (turingtest):

notice that makes our integrals quite tolerable :)

OpenStudy (anonymous):

\[Y_p(t)=-e^{-x}\int\frac{xe^{-x}e^{2x}}{e^{-2x}}dx+xe^{-x}\int\frac{e^{-x}e^{2x}}{e^{-2x}}dx\] That's a lot easier to integrate

OpenStudy (turingtest):

oh yeah :) especially after a little simplification

OpenStudy (anonymous):

\[-\frac{e^{2x}}{9}(3x-1)+\frac x3e^{3x}\]

OpenStudy (anonymous):

so what was the point of doing variation of parameters

OpenStudy (turingtest):

well now you have a particular solution

OpenStudy (turingtest):

y(x)=yc+yp you got you now so you're good to go

OpenStudy (turingtest):

you got yp*

OpenStudy (turingtest):

I did not check your integral btw

OpenStudy (anonymous):

that's ok I did wolfram

OpenStudy (turingtest):

right-o, I'd do the same right now :)

OpenStudy (anonymous):

oh I see \[c_1e^{-x}+c_2xe^{-x}-\frac{e^{2x}}{9}(3x-1)+\frac x3e^{3x}\]

OpenStudy (turingtest):

yep, and of course if this is an IVP we now proceed to find c1 and c2 otherwise we're done having found the general solution

OpenStudy (anonymous):

good, wow, that was a lot THanks Turing!

OpenStudy (turingtest):

Always a pleasure Sofiya :D

OpenStudy (mendicant_bias):

Looking back over this, I don't understand how you guys got a negative double root to the characteristic equation. I got m_1 = m_2 = +1. Plugging into the normal formula for repeated roots, getting e^x and xe^x, not e^-x. Am I missing something? @ganeshie8

OpenStudy (mendicant_bias):

@jim_thompson5910 @agent0smith

OpenStudy (mendicant_bias):

@Zarkon @terenzreignz Is anybody else familiar with ODE's seeing this, or am I missing something?

OpenStudy (mendicant_bias):

@zepdrix Anybody?

zepdrix (zepdrix):

@Mendicant_Bias For the homogeneous solution? Yes the characteristic equation appears to be giving +1 as a repeated root, as you indicated

zepdrix (zepdrix):

So yah, that was a mistake early in the thread :C